There are 7 black balls and 9 red balls in an urn. If 3 balls are drawn without replacement, what is the probability that exactly 1 black ball is drawn? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

probability that exactly 1 black ball is drawn =P(1 black ball from 7 black ball and 2 red from 9 red balls)

=7C1*9C2/16C3 =7*36/560 =63/140 =0.4500